# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ Appendix A Classified sources from the literature
The following table contains references and the number of sources (per reference) obtained from the literature. We have taken care to remove any duplicates in the catalog, as well as to provide the most recent and precise classifications. We also excluded known foreground sources but we kept any other source that was probably member of the corresponding galaxy.
@@
Table 7: List of references with their corresponding number of sources that contribute to our collected literature catalog.
Galaxy (Total) Reference Sources
WLM (87) Venn et al. (2003) 2
Bresolin et al. (2006) 38
Levesque & Massey (2012) 11
Britavskiy et al. (2015) 13
Britavskiy et al. (2019) 1
Maravelias et al. (2023) 1
Bonanos et al. (2024) 21
NGC 55 (279) Castro et al. (2008) 168
Patrick et al. (2017) 11
Maravelias et al. (2023) 1
Bonanos et al. (2024) 99
IC 10 (137) Massey et al. (2007b) 30
Massey et al. (2007a) 3
Tehrani et al. (2017) 29
Britavskiy et al. (2019) 6
de Wit et al. 2025 (subm.) 69
M31 (1170)* Massey et al. (2009) 2
Drout et al. (2009) 18
Neugent et al. (2012) 3
Massey et al. (2016) 951
Gordon et al. (2016) 81
Shara et al. (2016) 1
Humphreys et al. (2017) 6
Massey et al. (2019) 17
Neugent et al. (2019) 37
Kraus (2019) 11
Sholukhova et al. (2020) 2
Sarkisyan et al. (2022) 1
Neugent & Massey (2023) 40
NGC 247 (64) Solovyeva et al. (2020) 2
Maravelias et al. (2023) 1
Bonanos et al. (2024) 61
NGC 253 (81) Comerón et al. (2003a) 11
Heida et al. (2015) 1
Maravelias et al. (2023) 1
Bonanos et al. (2024) 78
NGC 300 (793) Bresolin et al. (2002) 662
Schild et al. (2003) 46
Crowther et al. (2007) 103
Gazak et al. (2015) 27
Roth et al. (2018) 504
González-Torà et al. (2022) 164
Maravelias et al. (2023) 2
Bonanos et al. (2024) 122
141414 * Using the source list from Paper I. Updated classifications for three candidate to confirmed LBVs (Sarkisyan et al. 2022; Sholukhova et al. 2020), and 11 WR stars (Neugent & Massey 2023). Removed two sources found as duplicates (original IDs: M31-957 and M31-982). Added WRs from Neugent & Massey (2023). 1 Spectral types are from Comerón et al. (2003a), while corrected coordinates from Comerón et al. (2003b). 2 Updated entries with the most recent $V$ magnitudes for 40 BSGs from Bresolin et al. (2005). 3 Updated the list of sources presented in Schild et al. (2003), by adding 10 more sources and removing 2 (as non-WR). 4 Updated spectral types for sources initially in Roth et al. (2018) (these entries were removed from the corresponding list).
Table 8: Continuation of the table.
IC 1613 (86) Kurtev et al. (2001) 1
Bresolin et al. (2007) 53
Herrero et al. (2012) 1
Garcia & Herrero (2013) 12
Britavskiy et al. (2019) 3
Chun et al. (2022) 105
de Wit et al. 2025 (subm.) 6
M33 (1547)** Massey et al. (1996) 3
Massey (1998) 52
Massey & Johnson (1998) 48 6
Urbaneja et al. (2002) 4
Bruhweiler et al. (2003) 2
Urbaneja et al. (2005) 11 7
Massey et al. (2007a) 12
Neugent & Massey (2011) 2
Drout et al. (2012) 11
Humphreys et al. (2014) 4
Massey et al. (2016) 1190
@@
Martin & Humphreys (2017) 2
Humphreys et al. (2017) 24
Kourniotis et al. (2018) 4
Massey et al. (2019) 10 6,8
Neugent et al. (2019) 46
Kraus (2019) 7
Maryeva et al. (2020) 2 9
Smith et al. (2020) 1 9
Neugent (2021) 82 10
Liu et al. (2022) 18
Phoenix Dwarf (8) Menzies et al. (2008) 2
Saviane et al. (2009) 1
Britavskiy et al. (2015) 5
NGC 1313 (97) Hadfield & Crowther (2007) 80
Bonanos et al. (2024) 17
NGC 2366 (0) n/a 0
NGC 2403 (68) Humphreys et al. (2019) 27
Bresolin et al. (2022) 40
Kaldybekova et al. (2023) 1
M81 (82) Kudritzki et al. (2012) 26
Khan et al. (2013) 7
Humphreys et al. (2019) 28
Gómez-González et al. (2020) 21
Sextans B (2) Britavskiy et al. (2019) 2
NGC 3109 (127) Evans et al. (2007) 90
Flores-Durán et al. (2017) 1711
Davidge (2018) 3
Maravelias et al. (2023) 1
Bonanos et al. (2024) 16
NGC 3077 (0) n/a 0
151515 5 Two sources (stars 4 and 12) were removed as they provided broader classifications rather than the types in Britavskiy et al. (2019) originally. ** Using IDs from Paper I.
@@
7 IDs OB 10-3 and UIT 136 refer to the same object. Kept the latest classification from that paper.
@@
Table 9: Continuation of the table.
Sextans A (143) Kaufer et al. (2004) 5
Britavskiy et al. (2014) 2
Britavskiy et al. (2015) 6
Camacho et al. (2016) 9
Garcia et al. (2019) 2
Lorenzo et al. (2022) 106
Bonanos et al. (2024) 13
NGC 4214 (0) n/a 0
NGC 4736 (3) Solovyeva et al. (2019) 2
Solovyeva et al. (2021) 112
NGC 4826 (0) n/a 0
M83 (241) Hadfield et al. (2005) 10513
Bresolin et al. (2016) 14
Della Bruna et al. (2022) 6613,14
Bonanos et al. (2024) 56
NGC 5253 (0) n/a 0
NGC 6822 (83) Massey (1998) 9
Bianchi et al. (2001) 1
Massey et al. (2007a) 2
Levesque & Massey (2012) 12
Patrick et al. (2015) 10
de Wit et al. 2025 (subm.) 49
Pegasus DIG (9) Massey et al. (2007a) 1
Britavskiy et al. (2015) 6
Britavskiy et al. (2019) 2
NGC7793 (166) Bibby & Crowther (2010) 7815
Khan et al. (2013) 3
Della Bruna et al. (2020) 5
Wofford et al. (2020) 36
Della Bruna et al. (2021) 5
Maravelias et al. (2023) 1
Bonanos et al. (2024) 38
@@
161616 12 Magnitude derived from Solovyeva et al. (2019). 13 Assigned WR to all sources but kept the original classification within ”[]”. A magnitude value for the HeII $\lambda$4685 filter is provided wherever available else a flag value of -99 is assigned. 14 Removed WR64 as duplicate to WR44, but WR35 and WR37 were kept as different (separation at 0.999″). 15 Removed all ”not WR” objects (20), but kept ”WN?” and ”WC?” sources as WR candidates.
Appendix B Comparing literature to predicted classes
@@
In this Appendix we provide a detailed description of the comparison results between the literature and the classifier predictions per galaxy (except for NGC 2366, NGC 3077, NGC 4214, NGC 4826, and NGC 5253, for which no sources were found). For each case we provide the number of sources found in the literature and the statistics in two distinct groups, one regarding the best (more reliable) results for sources that fulfilled the quality criteria as defined in Section 2.4, and another one for the whole set.
@@
Regarding the comparison between the predicted class and the literature spectral type, we followed this matching algorithm: i. a predicted ”RSG” was assumed to be compatible with RSG or candidates, sources with K or M types (even broader cases such as ’early/late K/M’), carbon stars (’C-star’,’carbon’), broader labels such as ’red’ and ’cool star’, ii. ”WR” was compatible with any WR spectral type (WC, WN, etc) including candidates, iii. ”YSG” was compatible with any YSG, F/G (including ’early/late F/G’ labels), or ’warm supergiant’, iv. ”BSG” was compatible with any O/B/A [^17] The inclusion of A-type stars in this class follows the training of the method, see Paper I. (and broader classes such as ’early/mid/late B’), OB or ’Hot Supergiant’, v. ”BeBR” was compatible with any source with a B[e] spectral type or candidate, vi. ”LBV” was compatible with known and candidate LBVs, vii. ”GAL” was compatible with any label indicative of a galaxy (e.g. ’galaxy’, ’quasar’, ’AGN’, ’QSO’), but we also considered them matching with carbon stars (pointing to elliptical galaxies where older stellar populations dominate), viii. broad spectral classifications from the literature such as ’blue’ or ’hot’, emission type stars (including labels as ’Halpha star’ and ’OBem’), Hii regions and planetary nebulae (’PN’) were considered compatible with any (predicted) class of ”WR”, ”BeBR”, ”LBV”, or ”BSG”, ix. any other very coarse or doubtful case (such as ’composite’, ’cluster’, ’star’, ’SNR’, ’neb’, ’symbiotic’, ’foreground’) were considered uncertain.
@@
Regarding the matches between our catalog and the literature source we considered 1″search radius. These matches are referred to as ’good matches’. We point here that the result of the process depends also on the accuracy of the coordinates taken from the literature.
@@
In Table 10 we provide the first few rows of the catalog that contains all sources with spectral classification from the literature. It includes all massive stars and candidates known, as well as another $\sim 330$ sources of point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, clusters). We provide our own ID (with a preceding ’lit-’ to differentiate from our main catalog), an IDlit that corresponds to the ID provided in the corresponding work (if present). We provide the coordinates and the spectral type. For reference, a magnitude value in a specific filter is provided, as derived from the original paper and in some cases corresponds to an indicative value (e.g. in variable sources such as in LBVs). When not available, a flagged value of -99 is given. Instead of the typical citation style we give the full bibliography code as provided by NASA’s Astrophysics Data System.
@@
Table 10: Compiled catalog of sources with spectral classifications from the literature, including all massive stars and candidates, as well as additional point sources.
ID IDlit R.A. Dec. Spectral Type Magnitude Filter Bibcode
@@
lit-WLM-1 [SC85b] 15 0.498125 --15.490805 A5 Ib 18.1 V 2003AJ….126.1326V
lit-WLM-2 [SC85b] 31 0.502625 --15.475055 A5 Ib 18.4 V 2003AJ….126.1326V
lit-WLM-3 A12 0.471833 --15.477694 B9 Ia 17.98 V 2006ApJ…648.1007B
lit-WLM-4 A11 0.499875 --15.472000 O9.7 Ia 18.4 V 2006ApJ…648.1007B
lit-WLM-5 A14 0.498250 --15.490638 A2 II 18.43 V 2006ApJ…648.1007B
lit-WLM-6 A9 0.488375 --15.455027 B1.5 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-7 A16 0.491250 --15.503722 A2 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-8 A19 $0$-15.521111 G2 I 18.62 V 2006ApJ…648.1007B
lit-WLM-9 B12 0.502667 --15.474972 A2 II 18.77 V 2006ApJ…648.1007B
lit-WLM-10 B13 0.473546 --15.474972 B1 Ia 18.92 V 2006ApJ…648.1007B
 [^18] This table is available in its entirety in the VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
B.1 WLM
@@
A total of 87 sources were collected from 7 works in the literature, for which we obtained good matches (i.e. within 1″ ) for 62 sources ($\sim 71\%$). We first checked the statistics for sources with the most reliable predictions, amounting to 30 sources (almost half of the sample). We obtained an exact classification for 86.7% (26 sources) and an erroneous classification for 13.3% (4 sources). Considering the whole sample (62 sources) we got 71.0% (44 sources) as exact matches, 4.8% (3) as matches within errors, 21.0% (13) as erroneous matches, and 3.2% (2) as uncertain. By combining the correct matches we have a 75.8% success rate.
B.2 NGC 55
@@
For NGC 55 we collected 279 sources from 4 different works. We removed 180 sources with no match at 1″  (equivalent to 64.5% of the sample). From the remaining 99 sources we did not have any reliable prediction according to the selection criteria we defined. Accounting for the whole sample we managed to obtain 14.1% (14 sources) as exact matches, 10.1% (10) as matches within errors, 73.7% (73) as erroneous matches, and 2.0% (2) as uncertain. The total success rate is 24.2%. We note here that almost all data in NGC 55 suffer from significant missing values (band completeness equal to 0.2), which means that the classifier cannot reach its top performance.
B.3 M31
As the closest and largest galaxy the number of spectroscopically confirmed objects we managed to collect summed to 1170 sources (combining 13 different works). A bit more than half of them (57.4%, 671 sources) do not have a good match. The majority of the remaining sources (382) passed the quality criteria with a 94.0% (359) being correctly predicted by the classifier, and only a 6.0% (23) classified erroneously. When considering all 499 sources the fraction of exact matches was 82.8% (413), with another 5.6% (28) predicted correctly within the error, and 11.6% (58) classified incorrectly. The total correct fraction is 88.4%. This high fractions (especially for the good candidates) are not unexpected since the classifier has been trained on (most) of these sources.
B.4 IC 10

### v2

@@ Appendix B Classified sources from the literature
The following table contains references and the number of sources (per reference) obtained from the literature. We have taken care to remove any duplicates in the catalog, as well as to provide the most recent and precise classifications. We also excluded known foreground sources, but we kept any other source that was probably a member of the corresponding galaxy.
@@
Table 7: References with their corresponding number of sources that contribute to our collected literature catalog. [^15] See notes in Appendix B.
Galaxy (Total) Reference Sources Galaxy (Total) Reference Sources
WLM (87) Venn et al. (2003) 2 M33 (1547)∗∗ Martin and Humphreys (2017) 2
Bresolin et al. (2006) 38 Humphreys et al. (2017) 24
Levesque and Massey (2012) 11 Kourniotis et al. (2018) 4
Britavskiy et al. (2015) 13 Massey et al. (2019) 106,8
Britavskiy et al. (2019) 1 Neugent et al. (2019) 46
Maravelias et al. (2023) 1 Kraus (2019) 7
Bonanos et al. (2024) 21 Maryeva et al. (2020) 29
NGC 55 (279) Castro et al. (2008) 168 Smith et al. (2020) 19
Patrick et al. (2017) 11 Neugent (2021) 8210
Maravelias et al. (2023) 1 Liu et al. (2022) 18
Bonanos et al. (2024) 99 Phoenix Dwarf (8) Menzies et al. (2008) 2
IC 10 (137) Massey et al. (2007b) 30 Saviane et al. (2009) 1
Massey et al. (2007a) 3 Britavskiy et al. (2015) 5
Tehrani et al. (2017) 29 NGC 1313 (97) Hadfield and Crowther (2007) 80
Britavskiy et al. (2019) 6 Bonanos et al. (2024) 17
de Wit et al. (2025) 69 NGC 2366 (0) n/a 0
M31 (1170) Massey et al. (2009) 2 NGC 2403 (68) Humphreys et al. (2019) 27
Drout et al. (2009) 18 Bresolin et al. (2022) 40
Neugent et al. (2012) 3 Kaldybekova et al. (2023) 1
Massey et al. (2016) 951 M81 (82) Kudritzki et al. (2012) 26
Gordon et al. (2016) 81 Khan et al. (2013) 7
Shara et al. (2016) 1 Humphreys et al. (2019) 28
Humphreys et al. (2017) 6 Gómez-González et al. (2020) 21
Massey et al. (2019) 17 Sextans B (2) Britavskiy et al. (2019) 2
Neugent et al. (2019) 37 NGC 3109 (127) Evans et al. (2007) 90
Kraus (2019) 11 Flores-Durán et al. (2017) 1711
Sholukhova et al. (2020) 2 Davidge (2018) 3
Sarkisyan et al. (2022) 1 Maravelias et al. (2023) 1
Neugent and Massey (2023) 40 Bonanos et al. (2024) 16
NGC 247 (64) Solovyeva et al. (2020) 2 NGC 3077 (0) n/a 0
Maravelias et al. (2023) 1 Sextans A (143) Kaufer et al. (2004) 5
Bonanos et al. (2024) 61 Britavskiy et al. (2014) 2
NGC 253 (81) Comerón et al. (2003a) 11 Britavskiy et al. (2015) 6
Heida et al. (2015) 1 Camacho et al. (2016) 9
Maravelias et al. (2023) 1 Garcia et al. (2019) 2
Bonanos et al. (2024) 78 Lorenzo et al. (2022) 106
NGC 300 (793) Bresolin et al. (2002) 662 Bonanos et al. (2024) 13
Schild et al. (2003) 46 NGC 4214 (0) n/a 0
Crowther et al. (2007) 103 NGC 4736 (3) Solovyeva et al. (2019) 2
Gazak et al. (2015) 27 Solovyeva et al. (2021) 112
Roth et al. (2018) 504 NGC 4826 (0) n/a 0
González-Torà et al. (2022) 164 M83 (241) Hadfield et al. (2005) 10513
Maravelias et al. (2023) 2 Bresolin et al. (2016) 14
Bonanos et al. (2024) 122 Della Bruna et al. (2022) 6613,14
IC 1613 (86) Kurtev et al. (2001) 1 Bonanos et al. (2024) 56
Bresolin et al. (2007) 53 NGC 5253 (0) n/a 0
Herrero et al. (2012) 1 NGC 6822 (83) Massey (1998) 9
Garcia and Herrero (2013) 12 Bianchi et al. (2001) 1
Britavskiy et al. (2019) 3 Massey et al. (2007a) 2
Chun et al. (2022) 105 Levesque and Massey (2012) 12
de Wit et al. (2025) 6 Patrick et al. (2015) 10
M33 (1547)∗∗ Massey et al. (1996) 3 de Wit et al. (2025) 49
Massey (1998) 52 Pegasus DIG (9) Massey et al. (2007a) 1
Massey and Johnson (1998) 486 Britavskiy et al. (2015) 6
Urbaneja et al. (2002) 4 Britavskiy et al. (2019) 2
Bruhweiler et al. (2003) 2 NGC 7793 (166) Bibby and Crowther (2010) 7815
Urbaneja et al. (2005) 117 Khan et al. (2013) 3
Massey et al. (2007a) 12 Della Bruna et al. (2020) 5
Neugent and Massey (2011) 2 Wofford et al. (2020) 36
Drout et al. (2012) 11 Della Bruna et al. (2021) 5
Humphreys et al. (2014) 4 Maravelias et al. (2023) 1
Massey et al. (2016) 1190 Bonanos et al. (2024) 38
@@
Notes to Table 7.
Using the source list from Paper I. Updated classifications for three candidate to confirmed LBVs (Sarkisyan et al. 2022; Sholukhova et al. 2020), and 11 WR stars (Neugent and Massey 2023). Removed two sources found as duplicates (original IDs: M31-957 and M31-982). Added WRs from Neugent and Massey (2023).
∗∗ Using IDs from Paper I.
1 Spectral types are from Comerón et al. (2003a), while corrected coordinates from Comerón et al. (2003b).
2 Updated entries with the most recent $V$ magnitudes for 40 BSGs from Bresolin et al. (2005).
3 Updated the list of sources presented in Schild et al. (2003), by adding 10 more sources and removing 2 (as non-WR).
4 Updated spectral types for sources initially in Roth et al. (2018) (these entries were removed from the corresponding list).
5 Two sources (stars 4 and 12) were removed as they provided broader classifications rather than the classification provided by the older work of Britavskiy et al. (2019).
@@
7 IDs OB 10-3 and UIT 136 refer to the same object. We kept the latest classification from that paper.
@@
12 Magnitude derived from Solovyeva et al. (2019).
@@
13 Assigned WR to all sources but kept the original classification within ”[]”. A magnitude value for the HeII $\lambda$4685 filter is provided wherever available; elsewhere a flag value of -99.0 is assigned.
14 Removed WR64 as duplicate to WR44, but WR35 and WR37 were kept as different (separation at 0.999″).
15 Removed all ”not WR” objects (20), but kept ”WN?” and ”WC?” sources as WR candidates.
Appendix C Comparing the literature to predicted classes
@@
In this appendix, we provide a detailed description of the comparison results between the literature and the classifier predictions per galaxy (except for NGC 2366, NGC 3077, NGC 4214, NGC 4826, and NGC 5253, for which no sources were found). For each case, we provide the number of sources found in the literature and the statistics in two distinct groups, one regarding the best (more reliable) results for sources that fulfilled the quality criteria as defined in Sect. 2.4, and another for the whole set.
@@
Regarding the comparison between the predicted class and the literature spectral type, we followed this matching algorithm: i. a predicted ”RSG” was assumed to be compatible with RSG or candidates, sources with K or M types (even broader cases such as ”early or late K or M”), carbon stars (”C-star”,”carbon”), broader labels such as ”red” and ”cool star”; ii. ”WR” was compatible with any WR spectral type (WC, WN, etc.) including candidates; iii. ”YSG” was compatible with any YSG, F or G (including ”early or late F or G” labels), or ”warm supergiant”; iv. ”BSG” was compatible with any of the O, B, or A-type stars [^16] The inclusion of A-type stars in this class follows the training of the method, see Paper I. (and broader classes such as ”early, mid, or late B”), OB or ”hot supergiant”; v. ”BeBR” was compatible with any source with a B[e] spectral type or candidate; vi. ”LBV” was compatible with known and candidate LBVs; vii. ”GAL” was compatible with any label indicative of a galaxy (e.g., ”galaxy”, ”quasar”, ”AGN”, and ”QSO”), but we also considered them matching with carbon stars (pointing to elliptical galaxies where older stellar populations dominate); viii. broad spectral classifications from the literature such as ”blue” or ”hot”, emission type stars (including labels as ”Halpha star” and ”OBem”), Hii regions and planetary nebulae (”PN”) were considered compatible with any (predicted) class of ”WR”, ”BeBR”, ”LBV”, or ”BSG”; ix. any other very coarse or doubtful case (such as ”composite”, ”cluster”, ”star”, ”SNR”, ”neb”, ”symbiotic”, ”foreground”) were considered uncertain.
@@
Regarding the matches between our catalog and the literature source, we considered a 1″search radius. These matches are referred to as ”good matches”. We point out here that the result of the process also depends on the accuracy of the coordinates taken from the literature.
From the total number of sources found in the literature for each galaxy, we first excluded those without matches within the search radius (as defined above). In addition, we also excluded ambiguous sources from the literature with very broad classifications (such as ”Cluster”, ”Composite”, ”Star”, ”SNR”, ”Nebula”). From the remaining sources, we calculated the success rate by combining the number of secure and candidate sources over the total number of available sources. We provide detailed numbers for each of this group of sources.
@@
In Table 8 we provide the first few rows of the catalog that contains all sources with spectral classification from the literature. It includes all known massive stars and candidates, as well as another $\sim 330$ sources of point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, and clusters). We provide our own ID (with a preceding ”lit-” to differentiate from our main catalog), an IDlit that corresponds to the ID provided in the corresponding work (if present). We provide the coordinates and the spectral type. For reference, a magnitude value in a specific filter is provided, as derived from the original paper, and in some cases corresponds to an indicative value (e.g., in variable sources such as in LBVs). When not available, a flagged value of -99.0 is given. Instead of the typical citation style, we give the full bibliography code as provided by NASA’s Astrophysics Data System.
@@
Table 8: Compiled catalog of sources with spectral classifications from literature, including all massive stars and candidates, as well as additional point sources.  [^17] This table is available in its entirety in VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
ID LitID RAJ2000 DEJ2000 SpecType Magnitude Filter Bibcode
@@
lit-WLM-1 [SC85b] 15 0.49812 -15.49081 A5 Ib 18.1 V 2003AJ….126.1326V
lit-WLM-2 [SC85b] 31 0.50262 -15.47506 A5 Ib 18.4 V 2003AJ….126.1326V
lit-WLM-3 A12 0.47183 -15.47769 B9 Ia 17.98 V 2006ApJ…648.1007B
lit-WLM-4 A11 0.49988 -15.47200 O9.7 Ia 18.4 V 2006ApJ…648.1007B
lit-WLM-5 A14 0.49825 -15.49064 A2 II 18.43 V 2006ApJ…648.1007B
lit-WLM-6 A9 0.48838 -15.45503 B1.5 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-7 A16 0.49125 -15.50372 A2 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-8 A19 0.50350 -15.52111 G2 I 18.62 V 2006ApJ…648.1007B
lit-WLM-9 B12 0.50267 -15.47497 A2 II 18.77 V 2006ApJ…648.1007B
lit-WLM-10 B13 0.47354 -15.47497 B1 Ia 18.92 V 2006ApJ…648.1007B
C.1 WLM
@@
A total of 87 sources were collected from seven works in the literature, for which we obtained good matches (i.e., within 1″ ) for 62 sources ($\sim 71\%$), resulting in 60 sources after removing two additional ambiguous literature classifications. We first checked the statistics for sources with the most reliable predictions, amounting to 30 sources (almost half of the sample). We obtained secure classifications for 86.7% (26 sources) and uncertain classifications for 13.3% (4 sources). Considering the whole sample (60 sources), we got 73.3% (44 sources) secure classifications, 5.0% (3) candidates, and 21.0% (13) as uncertain. By combining the secure classifications and the candidates, we obtain a success rate of 78.3%.
C.2 NGC 55
@@
For NGC 55 we collected 279 sources from four different works. We removed 180 sources with no match at 1″  (equivalent to 64.5% of the sample), and two other ambiguously classified objects. From the remaining 97 sources we did not have any reliable prediction according to the selection criteria we defined. Accounting for the whole sample, we managed to obtain 1.0% (1 source) as secure, 21.6% (21) as candidates, and 77.3% (73) as uncertain classifications. The total success rate is 22.7%. We note here that almost all data in NGC 55 suffer from significant missing values (band completeness equal to 0.2), which means that the classifier cannot reach its top performance.
C.3 IC 10